OpenStudy (anonymous):

Solve the equation: −4(x + 10) − 6 = −3(x − 2)

OpenStudy (igreen):

Distribute -4 and -3 into both of the parentheses: \(−4(x + 10) − 6 = −3(x − 2)\) \(−4x - 40 − 6 = −3x + 6\)

OpenStudy (geekfromthefutur):

-4(x + 10) - 6 = -3(x - 2) Multiply to clear parentheses -4x - 40 - 6 = -3x + 6 -4x - 46 = -3x + 6 Add 46 to both sides of the equation -4x - 46 + 46 = -3x + 6 + 46 -4x = -3x + 52 Add -3x to both sides of the equation -4x + 3x = -3x + 52 + 3x -x = 52 Multiply by -1 on both sides of the equation x =

OpenStudy (igreen):

x = -52

OpenStudy (geekfromthefutur):

Thats the answer its -52

OpenStudy (anonymous):

thanks!!!

OpenStudy (geekfromthefutur):

Anytime